OpenStudy (anonymous):

Evaluate: \[\sum_{i=0}^{\infty}(1/2)^{i}\]

OpenStudy (across):

2

OpenStudy (eyust707):

why is that again?

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

The formula for infinite geometric series is a/(1-r), where a is the starting value and r is the ratio. In this case, 1/2 is the ratio, and 1 is the starting value. 1/(1-(1/2)) = 1/(1/2) = 2.

OpenStudy (across):

EdgeOfTheEarth explained it better. :)

OpenStudy (eyust707):

suhweet

OpenStudy (eyust707):

sry for hijackin your thread marly =P